OpenStudy (anonymous):

Can someone help me find an equation for the inverse of the function. g(x)=x-3/2

OpenStudy (campbell_st):

is the equation \[g(x) = \frac{x -3}{2}\]

OpenStudy (anonymous):

To find the inverse of an equation, flip the x and y values first. Your equation will turn from \[g(x)=\frac{ x-3 }{ 2 }\] to \[x =\frac{ y -3 }{ 2 }\] After you've done that, solve for y.

OpenStudy (anonymous):

Yes @campbell_st

OpenStudy (anonymous):

Is y 1/2?

OpenStudy (anonymous):

To get the inverse, solve for y. By solving for y, you need to isolate it. \[x =\frac{ y -3 }{ 2 }\] Multiply both sides of the equation by 2 to get rid of the division sign.